100=20-5r
please solve.
100=20-5r
-20 -20
_______
80=-5r
answer: -16=r



Found 3 solutions by THANApHD, stanbon, minnie:
Answer by THANApHD(104)

-5r = 80
r = 80/-5
r = -16

Answer by stanbon(48535)

100=20-5r
Add 5r to both sides to get:
100+5r=20
Subtract 100 from both sides to get:
5r=20-100
5r=-80
Divide both sides by 5 to get:
r=-16
Cheers,
Stan H.

Answer by minnie(1)